Alex can complete a work in 5 hours and Ben can complete the same work in 10 hours .How many hours will they both take to complete the job together?
Found 2 solutions by checkley79, nerdybill:
Answer by checkley79(3341)

1/5+1/10=1/x
(2+1)/10=1/x
3/10=1/x
3x=10
x=10/3
x=3.33 hours.

Answer by nerdybill(7384)

Alex can complete a work in 5 hours and Ben can complete the same work in 10 hours .How many hours will they both take to complete the job together?
.
Let x = time (hours) for both
then
x(1/5 + 1/10) = 1
multiplying both sides by 10:
x(2 + 1) = 10
x(3) = 10
x = 10/3
x = 3 and 1/3 hours
or
x = 3 hours and 20 minutes